A vessel contained 1 mole of O2 gas (molar mass 32) at a temperature T. The pressure of the gas is P. An identical vessel containing one mole of the gas (molar mass 4) at a temperature 2T has a pressure of: (A) P/8 (B) P (C) 2 P (D) 8P 

Grade:12th pass

2 Answers

Riddhish Bhalodia
8 years ago
Assuming both the gases are ideal gases we can say that
PV/T = constant
and thus
PV/T = P’V/2T
=>
P’ = 2P which is option C
Yash kumavat
5 years ago
At constant V
P is directly proportional to T
So P/T = CONSTANT
So P1/T1=P2/T2
put P1=P .  T1=T .  T2=2T
      P2=?
P/T=P2/2T.  P2=2P
P2=2T×P/T
P2 =2P